# Off-shell supersymmetry via manifest invariance

### v1

@@ I Introduction
Finding a general off-shell formulation of supersymmetric field theories, a very desirable goal aiming towards their description as quantum field theories, is a major challenge of theoretical physics of fundamental interactions. For theories with more than eight supercharges such a formulation is typically problematic, both in the rigid supersymmetry (susy) and in the supergravity (sugra) cases. The difficulties are magnified when coupling $\mathcal{N}$-extended theories to hypermultiplets, i.e. matter multiplets with maximum spin 1/2 and whose number of d.o.f. remains stable when moving from massive to massless representations. An off-shell supersymmetric description in ordinary superspace $M^{D|\text{\tiny{N}}\mathcal{N}}$, with $\text{\tiny{N}}=2^{D/2}$ and $D$ the number of spacetime dimensions, can be formulated only for a very restricted class of supermultiplets [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. A general off-shell description of representations for $\mathcal{N}$-extended models, with fully manifest susy, was shown to be possible at the cost of including an infinite number of auxiliary fields, thereby circumventing the no-go theorems of susy stating that adding any finite number of auxiliary fields cannot help to achieve the closure of the off-shell algebra [11, 12, 13]. The most successful frameworks in which off-shell theories have been technically developed involving an infinite number of auxiliary fields are projective superspace [14, 15] and harmonic superspace [16]. However, the physical and group-theoretical meaning of having an infinite number of auxiliary fields remains unclear.
@@
The above two points should be emphasized separately, as in some cases it is possible to construct off-shell supermultiplets even with a finite number of fields, ensuring the matching of bosonic and fermionic degrees of freedom. However, the susy transformations do not close off-shell on the fields, they only close on-shell. Consequently, in such cases, one cannot speak of a manifestly off-shell supersymmetric theory. For example, as noted in [17] and more recently revisited in [18], this occurs in the case of the so-called double tensor multiplets [19, 20, 21]. Similarly in $\mathcal{N}$-extended models and in higher-dimensional theories, see e.g. [22]. Moreover, the opposite issue may arise, namely off-shell closure without matching of bosonic and fermionic degrees of freedom.
@@
The remaining of this work is thus organized as follows: In Section (II) we briefly recall the basics of the DFM and its application to the reduction of susy – while we refer the interested reader to the above cited literature on the DFM for further technical and conceptual details. In Section III we present our novel approach to off-shell susy invariance via the DFM. The procedure applies at both the kinematical and the dynamical levels. We also provide an example of application to an $\mathcal{N}$-extended model, that is $\mathcal{N}=2$, $D=4$ pure sugra in its geometric (a.k.a. rheonomic) formulation in superspace [31] – the dressing will be perturbative and invariance achieved at 1st order. Finally, in Section IV we discuss future developments and propose novel possible approaches to addressing point (i) using a finite number of fields.
@@ II.2 DFM and supersymmetry reduction
As far as standard susy (as opposed to unconventional susy [26]) is concerned, it was shown in [25] that the “gauge-fixing” conditions typically used to extract the d.o.f. of the Rarita-Schwinger (RS) spinor-vector and gravitino fields – e.g., 12 off-shell d.o.f. in $\mathcal{N}=1$, $D=4$ – are actually instances of the DFM. That is, solving the dressing functional constraints actually realises those fields as (self-dressed) relational variables. In both cases of rigid susy and sugra, what is commonly imposed is the so-called gamma-tracelessness condition on the fermionic 1-form $\psi^{\alpha}=\psi^{\alpha}_{\mu}dx^{\,\mu}$, namely $\gamma^{\mu}\psi_{\mu}=0$ (for simplicity, from now on we will omit all spinor indices). This condition is rewritten for the new dressed variable $\psi^{u}$, i.e. $\gamma^{\mu}\psi^{u}_{\mu}=0$, obtained by applying the DFM rule of thumb, and then explicitly solved for $u$, showing that it indeed qualifies as a dressing field after verifying how it transforms under susy. In this context, the dressing is field-dependent and non-local – and thus susy a substantive symmetry [34]. In the RS case, in $D$ spacetime dimensions, one gets
@@
where $\chi:=1/D\,\gamma^{\mu}\,{\psi}_{\mu}$ carries spin-1/2. The minimal coupling of the RS field with gravity is described via the Lorentz spin connection $\omega^{ab}$ and the soldering 1-form $e^{a}$ (the vierbein). It is obtained via covariantization,
@@ Acknowledgements.
J.F. is supported by the Austrian Science Fund (FWF), grant [P 36542], and by the Czech Science Foundation (GAČR), grant GA24-10887S. L.R. is supported by the research grant GrIFOS, funding scheme PNRR Young Researchers, MSCA Seal of Excellence (SoE), CUP E13C24003600006, ID SOE2024__\__0000103, of which this paper is part.

### v2

@@ I Introduction
Finding a general off-shell formulation of supersymmetric field theories, a very desirable goal aiming towards their description as quantum field theories, is a major challenge of theoretical physics of fundamental interactions. For theories with more than eight supercharges such a formulation is typically problematic, both in the rigid supersymmetry (susy) and in the supergravity (sugra) cases. The difficulties are magnified when coupling $\mathcal{N}$-extended theories to hypermultiplets, i.e. matter multiplets with maximum spin 1/2 and whose number of d.o.f. remains stable when moving from massive to massless representations. An off-shell supersymmetric description in ordinary superspace $M^{D|\text{\tiny{N}}\mathcal{N}}$, with $\text{\tiny{N}}=2^{D/2}$ and $D$ the number of spacetime dimensions, can be formulated only for a very restricted class of supermultiplets [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. A general off-shell description of representations for $\mathcal{N}$-extended models, with fully manifest susy, was shown to be possible at the cost of including an infinite number of auxiliary fields, thereby circumventing the no-go theorems of susy stating that adding any finite number of auxiliary fields cannot help to achieve the closure of the off-shell algebra [11, 12, 13]. The most successful frameworks in which off-shell theories have been technically developed involving an infinite number of auxiliary fields are projective superspace [14, 15] and harmonic superspace [16]. Another powerful approach to off-shell susy considers, instead, constrained “pure” spinors – see [17, 18] and references therein. However, the physical and group-theoretical meaning of having an infinite number of auxiliary fields or constrained fields remains unclear.
@@
The above two points should be emphasized separately, as in some cases it is possible to construct off-shell supermultiplets even with a finite number of fields, ensuring the matching of bosonic and fermionic degrees of freedom. However, the susy transformations do not close off-shell on the fields, they only close on-shell. Consequently, in such cases, one cannot speak of a manifestly off-shell supersymmetric theory. For example, as noted in [19] and more recently revisited in [20], this occurs in the case of the so-called double tensor multiplets [21, 22, 23]. Similarly in $\mathcal{N}$-extended models and in higher-dimensional theories, see e.g. [24].
@@
The remaining of this work is thus organized as follows: In Section (II) we briefly recall the basics of the DFM and its application to the reduction of susy – while we refer the interested reader to the above cited literature on the DFM for further technical and conceptual details. This more formal section provides the general mathematical template of the DFM, to better appreciate its field-theoretic application to the susy case. A more physics-oriented reader may start directly with Section III, where we present our novel approach to off-shell susy invariance via the DFM. The procedure applies at both the kinematical and the dynamical levels. We also provide an example of application to an $\mathcal{N}$-extended model, that is $\mathcal{N}=2$, $D=4$ pure sugra in its geometric (a.k.a. rheonomic) formulation in superspace [33] – the dressing will be perturbative and invariance achieved at 1st order. Finally, in Section IV we discuss future developments and propose novel possible approaches to addressing point (i) using a finite number of fields.
@@ II Dressing Field Method: basics
Let us observe that if a dressing is introduced by fiat, as additional d.o.f. not built from those of $\Phi$, i.e. it is a $\Phi$-independent dressing field, one is dealing with a new distinct theory: The dressing field is said ad hoc, the dressed fields cannot be interpreted as representing the physical relational content of the original theory, and the gauge symmetry of the new theory is said to be artificial [36, 37]. This encompasses as a special case the so-called Stueckelberg trick, whereby one implements a gauge symmetry in a theory via the introduction of extra d.o.f., the Stueckelberg fields, which are indeed ad hoc dressing fields. Looking at Stueckelberg models (like massive Yang-Mills and massive gravity) through the DFM lens, one would “undo” the result of the Stueckelberg trick, removing an artificial gauge symmetry.
@@ II.2 DFM and supersymmetry reduction
As far as standard susy (as opposed to unconventional susy [28]) is concerned, it was shown in [27] that the “gauge-fixing” conditions typically used to extract the d.o.f. of the Rarita-Schwinger (RS) spinor-vector and gravitino fields – e.g., 12 off-shell d.o.f. in $\mathcal{N}=1$, $D=4$ – are actually instances of the DFM. That is, solving the dressing functional constraints actually realises those fields as (self-dressed) relational variables. In both cases of rigid susy and sugra, what is commonly imposed is the so-called gamma-tracelessness condition on the fermionic 1-form $\psi^{\alpha}=\psi^{\alpha}_{\mu}dx^{\,\mu}$, namely $\gamma^{\mu}\psi_{\mu}=0$ (for simplicity, from now on we will omit all spinor indices). This condition is rewritten for the new dressed variable, i.e. $\gamma^{\mu}\psi^{u}_{\mu}=0$, obtained by applying the DFM rule of thumb, and then explicitly solved for $u$, showing that it indeed qualifies as a dressing field after verifying how it transforms under susy. In this context, the dressing is field-dependent and non-local – and thus susy a substantive symmetry [37]. In the RS case, in $D$ spacetime dimensions, we have that the RS field $\psi$ transforms under susy as follows:
$\psi\mapsto\psi+d\varepsilon,$ (6)
$\varepsilon$ being the susy parameter. Then, one introduces the (dressed) variable
$\psi^{u}:=\psi+du,\quad\text{that is}\quad\psi^{u}_{\mu}:=\psi_{\mu}+\partial_{\mu}u,$ (7)
and, solving the “dressing constraint” $\gamma^{\mu}\psi^{u}_{\mu}=0$ for $u$, one gets
@@
where $\chi:=1/D\,\gamma^{\mu}\,{\psi}_{\mu}$ carries spin-1/2. Notice that $\chi^{u}=0$ by construction. See [27] for details. The minimal coupling of the RS field with gravity is described via the Lorentz spin connection $\omega^{ab}$ and the soldering 1-form $e^{a}$ (the vierbein). It is obtained via covariantization,
@@ IV Discussion
Irrespective of the chosen formalism, the DFM offers a versatile framework for systematically reducing the symmetries of a theory, including susy. By properly building a dressing field (whenever possible), one can effectively reduce the symmetry, thus building manifestly relational, automatically invariant observables. This paves the way for a novel quantization strategy we call Relational Quantization, and which we elaborate in a forthcoming paper [58] (see also for preliminary results and examples [32]).
@@ Acknowledgements.
J.F. is supported by the Austrian Science Fund (FWF), grant [P 36542], and by the Czech Science Foundation (GAČR), grant GA24-10887S. L.R. is supported by the GrIFOS research project, funded by the Ministry of University and Research (MUR, Ministero dell’Università e della Ricerca, Italy), PNRR Young Researchers funding program, MSCA Seal of Excellence (SoE), CUP E13C24003600006, ID SOE2024__\__0000103, of which this paper is part.